Conditional Probability. Let's say, table below shows number of people who work for the company located in Manhattan and live in NY, NJ or CT.

Put any numbers in this table:

NY - NJ - CT

Male ... - ... - ...

Female ... - ... - ...

Based on data in your table find probability that randomly selected person from this company:

1) is Female AND lives in NY: P(F and NY);

2) is Male OR lives in NJ: P(M or NJ);

3) lives in CT IF person is Male: P(CT if M).
Answer #1

Answer:

The table below shows that number of people who work for the company located in Manhattan and live in NY, NJ or CT.

NY NJ CT Total
Male 20 40 16 76
Female 30 17 22 69
Total 50 57 38 145

The probability of random selected person is,

1) p(F and NY) = 69 / 145 + 50/145 = 0.82

2) P(M or NJ) = 76/145 + 57/145 = 0.92

3) P(CT if M) = 16/76 = 0.21